How much must be paid back on a loan of R3000 for six months if simple interest of 12% per annum is charged
1
Expert's answer
2021-09-21T05:03:21-0400

"I=P×R×T\\=3000×\frac{12}{100}×\frac{6}{12}\\=180\\Amount=principal + interest\\=3000+180\\=3180"